OpenStudy (pink.1d):

What are the explicit equation and domain for a geometric sequence with a first term of 2 and a second term of -8?

OpenStudy (pink.1d):

These are the choices: an = 2(-8)n - 1; all integers where n ≥ 1 an = 2(-8)n - 1; all integers where n ≥ 0 an = 2(-4)n - 1; all integers where n ≥ 0 an = 2(-4)n - 1; all integers where n ≥ 1

OpenStudy (pink.1d):

MEDAL AND FAN :)

OpenStudy (pink.1d):

@Vocaloid I know im annoying cuz i almost always ask u for help but ur always right and i actually learn stuff when u answer rather than those who just give me answers

Vocaloid (vocaloid):

@pink.1d you're not annoying, don't apologize just hold on a sec

OpenStudy (pink.1d):

I think its the third one but my friend said im wrong

OpenStudy (pink.1d):

okie:)

Vocaloid (vocaloid):

ahah, your friend was right, I'm sorry to say you're very close, though!

Vocaloid (vocaloid):

alright, let me show you why the 3rd answer isn't right if we start with n = 0, then the first term is 2(-4)(n) - 1, where n = 0. that means that the first term is -1, which doesn't match with the problem

OpenStudy (pink.1d):

oopps wow sometimes im so stupid

OpenStudy (pink.1d):

ok now i think it is

OpenStudy (pink.1d):

fourth one duh

Vocaloid (vocaloid):

yup

OpenStudy (pink.1d):

thanks for helping me when i became stupid:) lol

OpenStudy (pink.1d):

What are the explicit equation and domain for an arithmetic sequence with a first term of 7 and a second term of 3? (5 points) an = 7 - 4(n - 1); all integers where n ≥ 1 an = 7 - 4(n - 1); all integers where n ≥ 0 an = 7 - 3(n - 1); all integers where n ≥ 1 an = 5 - 3(n - 1); all integers where n ≥ 0

OpenStudy (pink.1d):

@Vocaloid i think its

OpenStudy (pink.1d):

first one

OpenStudy (pink.1d):

also i have to ask u something

OpenStudy (pink.1d):

since you are more experienced at life , is it bad to listen to music while u do school work?

OpenStudy (pink.1d):

my mom keeps saying i will be a stupidhead and the info will go in one ear and out the other

Vocaloid (vocaloid):

@pink.1d yes, your answer is correct Hm, as for the music question, it really depends on the person. Studies have shown that most people perform worse when music is playing in the background with the exception of classical music.

OpenStudy (pink.1d):

ya thats what i found out

OpenStudy (pink.1d):

if u were my tutor i wud be able to do amazezayn the alg2 exam im being forced to take

OpenStudy (pink.1d):

wait no amazing no zayn he left :((((((( i hate him

OpenStudy (pink.1d):

jk i could never hate zayn

OpenStudy (pink.1d):

srry tht was my 1d moment